# Splitting differential equations using Galois theory

### v1

@@ 3.2. Groupoids, splitting and short exact sequences
Let us say a word about the assumptions of these lemmas before continuing. First, the assumption that the short exact sequence is $A$-definably split cannot be weakened: a counterexample is given in Jin’s thesis [9, Example 5.29]. Here is a quick presentation of it: fix $\mathcal{U}\models\mathrm{DCF}_{0}$ a sufficiently saturated model, and $\mathcal{C}$ its field of constants. If $s$ is a differential transcendental, Jin considers the $\mathcal{C}$-internal types $p$ generic of $\delta^{-1}(s)$ and $q=\delta^{-1}(p)$. He shows that the short exact sequence obtained is isomorphic to an extension of $G_{a}(\mathcal{C})$ by some unipotent group, and that the map $\delta:q\rightarrow p$ does not almost-split. Any such extension definably splits, but to define a splitting of the short exact sequence of binding groups, one needs a fundamental system of solutions for $q$.
As for the necessity of (normal-$\bigstar$), we did not find a counterexample, but we expect there is one.
@@ 4.1. Preliminaries on logarithmic-differential pullbacks
Using the lemmas 3.15, 3.16 and 3.18, we can use information about the splitting of definable group extensions to obtain information on the splitting of types. In this section, we are interested in the specific case, already investigated in [10] and [5], of types obtained by pullback via the logarithmic derivative, in a differentially closed field of characteristic zero.
For the rest of this section, we work in $\mathrm{DCF}_{0}$, fix some sufficiently saturated $\mathcal{U}\models\mathrm{DCF}_{0}$ with its derivation $\delta$. Also fix some algebraically closed differential subfield $F<\mathcal{U}$.
We let $\mathcal{C}$ be its field of constants. It is a pure algebraically closed field, and is stably embedded. In particular, this implies that definable groups in $\mathcal{C}$ are exactly algebraic groups, and definable maps between them are exactly morphisms of algebraic groups (see [17, Chapter 7, Section 4] for the equivalence of definable groups and definable morphisms with algebraic groups and morphisms).
@@
which is a surjective morphism from $(\mathcal{U},\cdot)$ to $(\mathcal{U},+)$.
@@
Let $p\in S_{1}(F)$. We define its pullback under the logarithmic derivative $\log_{\delta}$ to be $q:=\log_{\delta}^{-1}(p)$, the type of any $u$ such that $\log_{\delta}(u)\models p$ and $u\not\in\operatorname{acl}(F,\log_{\delta}(u))$.
It is a known fact that this defines a unique complete type, and that each fiber $q_{\log_{\delta}(u)}$ is a strongly minimal $\mathcal{C}$-internal type (see [9, Proposition 5.3]).
In his thesis [9], Jin conjectures (Conjecture 5.4) the following, for such types $p$ and $q$:
Conjecture 4.2.
If $p$ is almost $\mathcal{C}$-internal, the following are equivalent:
@@
    $q$ is almost $\mathcal{C}$-internal
@@
    $\log_{\delta}:q\rightarrow p$ is almost split
@@
    there is an integer $k\neq 0$ such that for some $u\models q$, there are $w_{1},w_{2}$ such that:
@@
    • $w_{1}\in\operatorname{dcl}(F,\log_{\delta}u)$
    • $\log_{\delta}(w_{2})\in\operatorname{dcl}(F)$
It is condition (3) that he and Moosa call splitting in their subsequent article [10]. To avoid confusion with our terminology, we will (temporarily) call this condition product-splitting.
In his thesis, Jin proves the implications $(3)\Rightarrow(2)\Rightarrow(1)$ (Remark 5.5). In their article [10], Jin and Moosa completely answer this question in the case where $p$ is the generic type of an equation $\delta(x)=f(x)$, where $f\in F(x)$ and is $\mathcal{C}$-internal. Their answer depends on the following fact, which itself is a consequence, in $\mathrm{DCF}_{0}$, of [17, Chapter 3, Corollary 1.6]:
Fact 4.3.
Let $(G,S)$ be a definable (in $\mathrm{DCF}_{0}$), faithful, transitive group action on a strongly minimal set $S$. Then $(G,S)$ is definably isomorphic to either:
@@
    $G_{a}(\mathcal{C})\rtimes G_{m}(\mathcal{C})$, acting on $\mathcal{C}$ by affine transformation,
@@
Suppose that $p\in S_{1}(F)$ is the generic type of $\delta(x)=f(x)$, for some $f\in F(X)$, and is $\mathcal{C}$-internal. They prove:
Theorem A of [10].
If the binding group of $p$ is not of dimension $3$, then $\log_{\delta}^{-1}(p)$ is almost $\mathcal{C}$-internal if and only if it product-splits (which implies that it is almost-split).
When the dimension is $3$, they give an example showing that we can have almost internality without product splitting, thus disproving Jin’s conjecture. More precisely, they show:
Theorem 4.4 of [10].
There exist a differential field $F$ and a minimal type $p\in S(F)$ internal to the constants, with binding group $\mathrm{PSL}_{2}(\mathcal{C})$, such that $\log_{\delta}^{-1}(p)$ is $\mathcal{C}$-internal but $\log_{\delta}:\log_{\delta}^{-1}(p)\rightarrow p$ does not product-splits.
They do not show, in this case, that this also implies that the logarithmic derivative map does not almost split (in our terminology). More generally, it was left open whether almost splitting and product-splitting were equivalent. Note that since this question does not involve $\mathcal{C}$-internality of $\log_{\delta}^{-1}(p)$, we need not assume $p$ to be almost $\mathcal{C}$-internal. We will prove that splitting and product splitting are equivalent in general. The following lemma will take care of the case where the fibers of $\log_{\delta}$ are not weakly $\mathcal{C}$-orthogonal (and will be useful later):
@@
If $n\neq m$, this gives the result by picking $w_{1}=\frac{d_{m}}{b_{n}}$ and $w_{2}=\mu$. Otherwise, we see that $\frac{d_{m}}{b_{m}}=\mu$, for some non-zero $\mu\in\mathcal{C}$. In this case, using euclidian division, we find $h\in L[X]$ with $\deg(h)<\deg(f)$ and such that $\frac{h(a)}{g(a)}=e-\mu$. As $e\not\in L$ and $\mu\in L$, we get $e-\mu\not\in L$. Since $\mu\in\mathcal{C}$ and $e\in\mathcal{C}$, we also get $e-\mu\in\mathcal{C}$. We can then repeat the proof using these new polynomials to conclude as before.
We now prove the equivalence of almost splitting and product-splitting:
@@
Let $q\in S_{1}(F)$ be a type over any algebraically closed differential subfield of $\mathcal{U}$. Let $p:=\log_{\delta}^{-1}(q)$. Then $\log_{\delta}:p\to q$ is almost split if and only if there is an integer $k\neq 0$ such that for some $a\models p$, there are $w_{1},w_{2}$ such that:
@@
As the types $s|_{L}$ and $p_{\log_{\delta}(a)}$ are interalgebraic, their binding groups must be isogenous by Corollary 2.13. By Lemma 4.4, we may assume that $p_{\log_{\delta}(a)}$ is weakly orthogonal to $\mathcal{C}$ and do so for the rest of the proof. Hence the binding group of $p_{\log_{\delta}(a)}$ is definably isomorphic to $G_{m}(\mathcal{C})$, and the binding group of $s|_{L}$ must be as well.
By Fact 2.5, the binding group $\mathrm{Aut}_{L}(s|_{L}/\mathcal{C})$ is a definable normal subgroup of $\mathrm{Aut}_{F}(s/\mathcal{C})$. Note that $\mathrm{RM}(s)=\mathrm{RM}(s|_{L})=\mathrm{RM}(p_{\log_{\delta}(a)})=1$, and as $F$ is algebraically closed, the type $s$ is strongly minimal. Since $s|_{L}$ is weakly $\mathcal{C}$-orthogonal forking calculus along with an automorphism argument yields that $s$ is also weakly $\mathcal{C}$-orthogonal. Hence its binding group, $\mathrm{Aut}_{F}(s/\mathcal{C})$, acts transitively (and faithfully) on the set of realizations of $s$. We apply Fact 4.3 and see that the only possible case is $\mathrm{Aut}_{F}(s/\mathcal{C})$ being definably isomorphic to $G_{m}(\mathcal{C})$.

### v3

@@ 3.2. Groupoids, splitting and short exact sequences
Let us say a word about the assumptions of these lemmas before continuing. First, the assumption that the short exact sequence is $A$-definably split cannot be weakened: a counterexample is given in Jin’s thesis [11, Example 5.29]. Here is a quick presentation of it: fix $\mathcal{U}\models\mathrm{DCF}_{0}$ a sufficiently saturated model, and $\mathcal{C}$ its field of constants. If $s$ is a differential transcendental, Jin considers the $\mathcal{C}$-internal types $p$ generic of $\delta^{-1}(s)$ and $q=\delta^{-1}(p)$. He shows that the short exact sequence obtained is isomorphic to an extension of $G_{a}(\mathcal{C})$ by some unipotent group, and that the map $\delta:q\rightarrow p$ does not almost-split. Any such extension definably splits, but to define a splitting of the short exact sequence of binding groups, one needs a fundamental system of solutions for $q$.
As for the necessity of ($\bigstar$), we did not find a counterexample, but we expect there is one.
Before we end this section, we include a result about a degenerated case of splitting, which will be of use later:
Lemma 3.17.
Let $p\in S(A)$ be a $\mathcal{C}$-internal type and $f:p\rightarrow f(p)$ be an $A$-definable fibration. Suppose that $f(p)$ is weakly $\mathcal{C}$-orthogonal and:
  • the short exact sequence $1\rightarrow K\rightarrow\mathrm{Aut}_{A}(p/\mathcal{C})\xrightarrow{\widetilde{f}}\mathrm{Aut}_{A}(f(p)/\mathcal{C})\rightarrow 1$ is $A$-definably split via some $A$-definable section $s$ and $K$ is finite,
  • the action of $\mathrm{Aut}_{A}(f(p)/\mathcal{C})$ satisfies ($\bigstar$).
Then the groupoid $\mathrm{Gd}_{A}(p,f/\mathcal{C})$ is retractable, thus by Lemma 3.11 we get a map $\pi:p\rightarrow\pi(p)$ such that $p$ is interdefinable with $\pi(p)\times f(p)$. Moreover $\pi(p)(\mathcal{U})\subset\operatorname{acl}(\mathcal{C},A)$.
Proof.
By Lemma 3.14, we know that the groupoid $\mathrm{Gd}_{A}(p,f/\mathcal{C})$ is retractable by the map $r$ defined by $r(f(a),f(b))=s(\sigma)|_{p_{f(a)}}$, for some (any) $\sigma\in\mathrm{Aut}_{A}(f(p)/\mathcal{C})$ with $\sigma(f(a))=f(b)$.
The map $\pi$ is defined to be the quotient map of the equivalence relation $E$ given by $E(x,y)$ if and only if $r(f(a),f(b))(a)=b$, for any $a,b\models p$. By the previous paragraph, this is equivalent to $s(\sigma)|_{p_{f(a)}}(a)=b$.
Let $\sigma\in\mathrm{Aut}_{A}(\mathcal{U}/\mathcal{C})$, and $a\models p$. Then there is $\tau\in K$ such that $\sigma|_{p}=s(\sigma|_{f(p)})\circ\tau$. We then have:
$\sigma(a)$ $=s(\sigma|_{f(p)})(\tau(a))$
$=s(\sigma|_{f(p)})|_{p_{f(\tau(a))}}(\tau(a))$
$=r\Bigl{(}f\bigl{(}\tau(a)\bigr{)},f\bigl{(}\sigma(\tau(a))\bigr{)}\Bigr{)}(\tau(a))$
$=r\Bigl{(}f\bigl{(}\tau(a)\bigr{)},f\bigl{(}\sigma(a)\bigr{)}\Bigr{)}(\tau(a))\text{ as }\tau\in K$
which implies that $E(\tau(a),\sigma(a))$. As $K$ is finite, this means that the orbit of $\pi(a)$ under $\mathrm{Aut}_{A}(\mathcal{U}/\mathcal{C})$ is finite. Hence $\pi(p)(\mathcal{U})\subset\operatorname{acl}(\mathcal{C},A)$. ∎
Before we move on to the specific case of $f$ being the logarithmic derivative in $\mathrm{DCF}_{0}$, we now state and prove the strongest result we could obtain in an $\omega$-stable context, which is Theorem A of the introduction:
Theorem 3.18.
Assume that $\mathcal{C}$ is a pure $A$-definable algebraically closed field. Let $p\in S(A)$ be a weakly $\mathcal{C}$-orthogonal type and $f:p\rightarrow q$ be an $A$-definable fibration. Assume that $q$ is $\mathcal{C}$-internal with a linear nilpotent binding group and that for any $a\models p$, the fiber $p_{f(a)}$ is $\mathcal{C}$-internal with diagonalizable binding group. Then $p$ is almost $\mathcal{C}$-internal if and only if $f$ almost splits.
Proof.
The right to left direction is immediate. For the left to right, we first assume that $p$ is outright $\mathcal{C}$-internal. By [8, Corollary 2.5], the binding group $\mathrm{Aut}_{A}(p/\mathcal{C})$ is connected. There is an $A$-definable short exact sequence:
$1\rightarrow K\rightarrow\mathrm{Aut}_{A}(p/\mathcal{C})\xrightarrow{\widetilde{f}}\mathrm{Aut}_{A}(q/\mathcal{C})\rightarrow 1$
By [8, Lemma 2.7] $K$ is definably isomorphic to $\mathrm{Aut}_{Af(\bar{a})}(\operatorname{tp}(\bar{a}/Af(\bar{a}))/\mathcal{C})$, where $\bar{a}=a_{1},\cdots,a_{n}$ is a Morley sequence in $p$ such that $\bar{a}$ is a fundamental system of realizations of $p$, and $f(\bar{a})$ is a fundamental system of realizations of $q$. As $\mathrm{Aut}_{Af(a)}(p_{f(a)}/\mathcal{C})$ is diagonalizable for any $f(a)\models q$, we see that $K$ is diagonalizable as a definable subgroup of the diagonalizable group $\prod\limits_{i=1}^{n}\mathrm{Aut}_{Af(a_{i})}(\operatorname{tp}(a_{i}/Af(a_{i}))/\mathcal{C})$. We would like to apply Lemma 2.18 to get an $A$-definable splitting of $f$. However, the group $K$ need not be connected in general, and we now show how $K$ can be replaced by its connected component by replacing $p$ with an interalgebraic type.
By Lemma 2.18, the binding group $\mathrm{Aut}_{A}(p/\mathcal{C})$ is definable linear nilpotent. In particular, the subgroup $K$ is central, as a subgroup of its maximal torus (which exists as $\mathrm{Aut}_{A}(p/\mathcal{C})$ is connected). Modifying the proof of [8, Lemma 2.2], we get that $K$ is $A$-definably isomorphic to the $\mathcal{C}$-points of a diagonalizable group $\widehat{K}$, defined over $A\cap\mathcal{C}$. Since $A\cap\mathcal{C}$ is an algebraically closed field, the diagonalizable group $\widehat{K}$ is split over $A\cap\mathcal{C}$ by [1, 8.11], meaning it is $A\cap\mathcal{C}$ definably isomorphic to an algebraic subgroup of $G_{m}(\mathcal{C})^{k}$ for some $k$. By [1, 8.7], it is $A\cap\mathcal{C}$-definably isomorphic to $\widehat{R}\times G_{m}(\mathcal{C})^{l}$, for some $l\geq 0$ and some finite, $A\cap C$-definable group $\widehat{R}$. Therefore $K$ is $A$-definably isomorphic to $R\times K^{\circ}$, for some $A$-definable finite group $R$.
The subgroup $R$ induces an $A$-definable equivalence relation $E$ on $p$ by $E(x,y)$ if and only if there is $\sigma\in R$ such that $\sigma(x)=y$. Let $\eta$ be the quotient map, it induces an $A$-definable group morphism $\widetilde{\eta}:\mathrm{Aut}_{A}(p/\mathcal{C})\rightarrow\mathrm{Aut}_{A}(\eta(p)/\mathcal{C})$.
Claim.
$\ker{\widetilde{\eta}}=R$.
Proof.
The kernel of $\widetilde{\eta}$ is the group of $\tau\in\mathrm{Aut}_{A}(p/\mathcal{C})$ such that for all $x\models p$, there is $\sigma\in R$ such that $\tau(x)=\sigma(x)$. By Lemma 2.19, for any $x\models p$, if $\tau$ is in the unipotent radical of $\mathrm{Aut}_{A}(p/\mathcal{C})$ and $\sigma$ is in its maximal torus, then $\tau(x)=\sigma(x)$ implies $\tau(x)=x=\sigma(x)$. As $R$ is a subgroup of the maximal torus and the action of $\mathrm{Aut}_{A}(p/\mathcal{C})$ on $p$ is faithful, Lemma 2.19 implies that $\ker(\widetilde{\eta})$ intersects the unipotent radical of $\mathrm{Aut}_{A}(p/\mathcal{C})$ trivially. Again by Lemma 2.19 (or rather, its proof), we see that $\ker(\widetilde{\eta})$ is a subgroup of the maximal torus of $\mathrm{Aut}_{A}(p/\mathcal{C})$. Now, if $\tau\in\ker(\widetilde{\eta})$, there must be some $x\models p$ and $\sigma\in R$ such that $\sigma(x)=\tau(x)$. Notice that the maximal torus is central in $\mathrm{Aut}_{A}(p/\mathcal{C})$, and $\mathrm{Aut}_{A}(p/\mathcal{C})$ acts transitively on $p$. From this, we deduce that the maximal torus acts freely on $p$, and therefore that $\sigma=\tau$. ∎
Since $R<\ker(\widetilde{f})$, we obtain a definable map $g:\eta(p)\rightarrow f(p)$ by letting $g(\eta(x))=f(x)$. Since $\eta$ has finite fiber, if we can show that $g:\eta(p)\rightarrow f(p)$ splits, then we automatically obtain almost splitting of $f:p\rightarrow f(p)$.
The map $g$ induces an $A$-definable short exact sequence:
$1\rightarrow K^{\circ}\rightarrow\mathrm{Aut}_{A}(\eta(p)/\mathcal{C})\xrightarrow{\widetilde{g}}\mathrm{Aut}_{A}(f(p)/\mathcal{C})\rightarrow 1$
The group $\mathrm{Aut}_{A}(\eta(p)/\mathcal{C})$ is, by Lemma 2.18, a definable linear nilpotent group, i.e. definably isomorphic to the $\mathcal{C}$-points of a linear nilpotent group. Moreover, the kernel $K^{\circ}$ is connected, so we can apply the second part of Lemma 2.18. We only need to show that the maximal tori of $\mathrm{Aut}_{A}(\eta(p)/\mathcal{C})$ and $\mathrm{Aut}_{A}(f(p)/\mathcal{C})$ are $A$-definably isomorphic to the $\mathcal{C}$-points of tori. Again, this can be obtained by noticing that they are central, and modifying the proof of [8, Lemma 2.2].
By Lemma 2.18, the short exact sequence induced by $g$ splits $A$-definably. We know that $G=\mathrm{Aut}_{A}(f(p)/\mathcal{C})=G_{u}\times T$, where $G_{u}$ is it unipotent radical and $T$ its maximal torus. By Lemma 2.19, for any $b\models f(p)$, we have $G_{u}b\cap Tb=\{b\}$. Because $G_{u}$ is definably unipotent, by Corollary 3.15, its action on $f(p)$ satisfies ()($\bigstar$)( ★ ). Again, since $\mathrm{Aut}_{A}(f(p)/\mathcal{C})$ acts transitively on $f(p)$ and $T$ is central in it, we see that the action of $T$ on $f(p)$ is free, and therefore satisfies ()($\bigstar$)( ★ ). By Lemma 3.16 the action of $\mathrm{Aut}_{A}(f(p)/\mathcal{C})$ on $f(p)$ satisfies ()($\bigstar$)( ★ ), and therefore by Lemma 3.14, the map $g:\eta(p)\rightarrow f(p)$ splits.
We now explain how to reduce the case where $p$ is almost $\mathcal{C}$-internal to the $\mathcal{C}$-internal case, using an argument pointed out to us by the anonymous referee. Consider the map $A$-definable map $g:p\rightarrow\widetilde{p}$ to from $p$ its largest $\mathcal{C}$-internal quotient, which is defined as follows: if $a\models p$, then $g(a)$ satisfies that
[…]
and in particular $\operatorname{tp}(g(a)/A)$ is $\mathcal{C}$-internal. A standard argument shows that this map exists. Moreover, it must have finite fibers because $p$ is almost $\mathcal{C}$-internal. As $q$ is $\mathcal{C}$-internal, we obtain a map $\widetilde{f}:\widetilde{p}\rightarrow q$ such that $f=\widetilde{f}\circ g$. Since the fibers of $g$ are finite, almost splitting of $f$ and $\widetilde{f}$ are equivalent.
The map $g$ induces, for any $a\models p$, a map on the fibers $p_{f(a)}\rightarrow\operatorname{tp}(g(a)/\widetilde{f}(a)A)$, and a routine argument shows that this induces a surjective map from the binding group of $p_{f(a)}$ to the binding group of $\operatorname{tp}(g(a)/\widetilde{f}(a)A)$. Therefore the binding group of $\operatorname{tp}(g(a)/\widetilde{f}(a)A)$ is diagonalizable as a quotient of a diagonalizable group. We apply the $\mathcal{C}$-internal case of the theorem to show that the map $\widetilde{f}$ is almost split, which implies that the map $f$ is also almost split. ∎
@@ 4.1. Preliminaries on logarithmic-differential pullbacks
Using lemmas 3.12, 3.14 and 3.16, we can use information about the splitting of definable group extensions to obtain information on the splitting of types. In this section, we are interested in the specific case, already investigated in [12] and [7], of types obtained by pullback via the logarithmic derivative, in a differentially closed field of characteristic zero.
For the rest of this section, we work with differentially closed fields of characteristic zero and assume familiarity with model theory of $\mathrm{DCF}_{0}$, for which a reference is [19]. Fix some sufficiently saturated $\mathcal{U}\models\mathrm{DCF}_{0}$ with its derivation $\delta$. Also fix some algebraically closed differential subfield $F<\mathcal{U}$.
We let $\mathcal{C}$ be the field of constants of $\mathcal{U}$. It is a pure algebraically closed field, and is stably embedded. In particular, this implies that definable groups in $\mathcal{C}$ are exactly algebraic groups, and definable maps between them are exactly morphisms of algebraic groups (see [19, Chapter 7, Section 4] for the equivalence of definable groups and definable morphisms with algebraic groups and morphisms).
@@
which is a surjective morphism from $(\mathcal{U}^{*},\cdot)$ to $(\mathcal{U},+)$ with kernel $(\mathcal{C}^{*},\cdot)$.
@@
Let $q\in S_{1}(F)$. We define its pullback under the logarithmic derivative $\log_{\delta}$ to be $p:=\log_{\delta}^{-1}(q)$, the type of any $u$ such that $\log_{\delta}(u)\models q$ and $u\not\in\operatorname{acl}(F,\log_{\delta}(u))$.
It is a known fact that this defines a unique complete type, and that each fiber $p_{\log_{\delta}(u)}$ is a strongly minimal $\mathcal{C}$-internal type (see [11, Proposition 5.3]).
Consider the following three conditions on $q$, the last two of which were introduced and studied in Jin’s thesis [11]:
@@
    $p=\log_{\delta}^{-1}(q)$ is almost $\mathcal{C}$-internal,
@@
    $\log_{\delta}:p\rightarrow q$ is almost split,
@@
    there is an integer $k\neq 0$ such that for some $u\models p$, there are $w_{1},w_{2}$ such that:
@@
    • $w_{1}\in\operatorname{dcl}(F,\log_{\delta}(u))$
    • $\log_{\delta}(w_{2})\in F$.
We will call condition $(3)$ product-splitting.
Jin proves the implications $(3)\Rightarrow(2)\Rightarrow(1)$ in [11, Remark 5.5]. In a subsequent article [12], Jin and Moosa consider the case where $q$ is $\mathcal{C}$-internal and the generic type of an equation $\delta(x)=f(x)$, for some $f\in F(x)$ (note that it is condition $(3)$ that they call splitting). Key to their work is the following result, itself a consequence, in $\mathrm{DCF}_{0}$, of [19, Chapter 3, Corollary 1.6]:
Fact 4.2.
Let $(G,S)$ be a definable (in $\mathrm{DCF}_{0}$), faithful, transitive action of a connected group $G$ on a strongly minimal set $S$. Then $(G,S)$ is definably isomorphic to either:
@@
    $G_{a}(\mathcal{C})\rtimes G_{m}(\mathcal{C})$, acting on $\mathcal{C}$ by affine transformations,
@@
They prove that if the binding group of $q$ is not isomorphic to $\mathrm{PSL}_{2}(\mathcal{C})$, then almost internality and product splitting are equivalent, and give an almost internal, not product-split example when it is isomorphic to $\mathrm{PSL}_{2}(\mathcal{C})$.
Jin and Moosa almost exclusively work with product-splitting, which has the advantage of having convenient algebraic characterizations (see Lemma 4.8). However, splitting is more convenient to work with when using binding groups, and is key to using the methods of Section 3. It is therefore desirable to understand the connection between splitting and product splitting. In what follows, we will prove, under various additional assumptions, the equivalence of product-splitting and almost splitting, without any dimension restriction.
The following lemma will be key in applications, and also shows that splitting of some fibrations can be reduced to splitting of the logarithmic derivative.
Lemma 4.3.
Let $F$ be an algebraically closed field, let $p\in S(F)$ be an almost $\mathcal{C}$-internal type and $f:p\rightarrow f(p)=q$ be a fibration, with $q$ being $\mathcal{C}$-internal and weakly $\mathcal{C}$-orthogonal. Suppose that for some (any) $a\models p$:
  • $p_{f(a)}$ is $\mathcal{C}$-internal and weakly $\mathcal{C}$-orthogonal,
  • $\mathrm{Aut}_{f(a)F}(p_{f(a)}/\mathcal{C})$ is definably isomorphic to $G_{m}(\mathcal{C})$.
Then there exists a type $p^{\prime}$, internal to $\mathcal{C}$, and a fibration $\log_{\delta}:p^{\prime}\rightarrow\log_{\delta}(p^{\prime})$ with $\log_{\delta}(p^{\prime})\subset\operatorname{dcl}(q(\mathcal{U})F)$ and the property that if $\log_{\delta}:p^{\prime}\rightarrow\log_{\delta}(p^{\prime})$ almost splits then $f:p\rightarrow f(p)$ almost splits.
Proof.
By Fact 3.6, there is a finite-to-one definable map $\pi:p\rightarrow\pi(p)$ such that $\pi(p)$ is $\mathcal{C}$-internal. In particular, for any $a\models p$, it restricts to a finite-to-one map from $p_{f(a)}$ to $\operatorname{tp}(\pi(a)/f(a)F)$. By Lemma 2.10, there is a finite-to-one definable morphism $\mathrm{Aut}_{f(a)F}(p_{f(a)}/\mathcal{C})\rightarrow\mathrm{Aut}_{f(a)F}(\operatorname{tp}(\pi(a)/f(a)F)/\mathcal{C})$. As the former group is definably isomorphic to $G_{m}(\mathcal{C})$, so is the latter.
Fix some $a\models p$. Since $p_{f(a)}$ is weakly $\mathcal{C}$-orthogonal, so is $\operatorname{tp}(\pi(a)/f(a)F)$. Moreover, as $q=f(p)$ is weakly orthogonal to $\mathcal{C}$, we have that $\mathcal{C}\cap F\langle f(a)\rangle=\mathcal{C}\cap F$, which is algebraically closed. By [12, Lemma 4.1], there is $b\in F\langle f(a),\pi(a)\rangle\setminus F\langle f(a)\rangle^{\mathrm{alg}}$ such that $\log_{\delta}(b)\in F\langle f(a)\rangle$. We let $p^{\prime}=\operatorname{tp}(b/F)$ and obtain a fibration $\log_{\delta}:p^{\prime}\rightarrow\log_{\delta}(p^{\prime})$ with $\mathcal{C}$-internal strongly minimal fibers. Finally as $b\in\operatorname{dcl}(f(a),\pi(a),F)$, we see that $p^{\prime}$ is $\mathcal{C}$-internal.
[…]
Claim.
We have $U(c/F)=1$.
Proof of claim.
[…]
$=U(c\log_{\delta}(b)/F)-U(\log_{\delta}(b)/F)$
$=U(b/F)-U(b/F)+1$
$=1$
[…]
By construction $b\in\operatorname{dcl}(aF)$, and $c\in\operatorname{acl}(bF)$, so $(c,f(a))\in\operatorname{acl}(aF)$. Note that $U(c,f(a)/F)=U(c/F)+U(f(a)/F)=1+U(f(a)/F)$. Since $\mathrm{Aut}_{f(a)F}(p_{f(a)}/\mathcal{C})$ is definably isomorphic to $G_{m}(\mathcal{C})$ and $p_{f(a)}$ is weakly $\mathcal{C}$-orthogonal, we see that $U(p_{f(a)})=1$, which implies that $U(f(a)/F)=U(a/F)-1$, and combining this with the previous equality we get $U(c,f(a)/F)=U(a/F)$. As $(c,f(a))\in\operatorname{acl}(aF)$, this implies that $a$ and $(c,f(a))$ are interalgebraic over $F$, and $f:p\rightarrow f(p)$ splits. ∎
We now start our study of pullbacks under the logarithmic derivative. The following lemma will take care of the case where the fibers of $\log_{\delta}$ are not weakly $\mathcal{C}$-orthogonal (and will be useful later):
@@
If $n\neq m$, this gives the result by picking $w_{1}=\frac{d_{m}}{b_{n}}$ and $w_{2}=\mu$. Otherwise, we see that $\frac{d_{m}}{b_{m}}=\mu$, for some non-zero $\mu\in\mathcal{C}$. In this case, using euclidian division, we find $h\in L[X]$ with $\deg(h)<\deg(f)$ and such that $\frac{h(a)}{g(a)}=e-\mu$. As $e\not\in L$ and $\mu\in L$, we get $e-\mu\not\in L$. Since $\mu\in\mathcal{C}$ and $e\in\mathcal{C}$, we also get $e-\mu\in\mathcal{C}$. We can then repeat the proof using these new polynomials to conclude as before. ∎
We now prove the equivalence of almost splitting and product-splitting, as long as the fibration is not degenerate in a precise sense:
@@
Let $q\in S_{1}(F)$ be a type over any algebraically closed differential subfield of $\mathcal{U}$. Let $p:=\log_{\delta}^{-1}(q)$. Suppose that $p(\mathcal{U})\not\subset\operatorname{acl}(\log_{\delta}(p)(\mathcal{U}),\mathcal{C},F)$. Then the fibration $\log_{\delta}:p\to q$ is almost split if and only if it is product-split, i.e. there is an integer $k\neq 0$ such that for some $a\models p$, there are $w_{1},w_{2}$ such that:
@@
As stated previously, that product-splitting implies almost splitting is essentially contained in [11, Remark 5.5].
@@
As the types $s|_{L}$ and $p_{\log_{\delta}(a)}$ are interalgebraic, their binding groups must be isogenous by Corollary 2.13. By Lemma 4.4, we may assume that $p_{\log_{\delta}(a)}$ is weakly orthogonal to $\mathcal{C}$ and do so for the rest of the proof. Hence the binding group of $p_{\log_{\delta}(a)}$ is definably isomorphic to $G_{m}(\mathcal{C})$, and the binding group of $s|_{L}$ must be as well. We want to show that $\mathrm{Aut}_{F}(s/\mathcal{C})$ is also definably isomorphic to $G_{m}(\mathcal{C})$.
First, remark that $\mathrm{Aut}_{F}(s/\mathcal{C},\log_{\delta}(p))$ is infinite. Suppose, on the contrary, that it is finite. Then $s(\mathcal{U})\subset\operatorname{acl}(\mathcal{C},\log_{\delta}(p)(\mathcal{U}),F)$, and as $p$ is interalgebraic with $s\otimes\log_{\delta}(p)$ over $F$, this would imply that $p(\mathcal{U})\subset\operatorname{acl}(\mathcal{C},\log_{\delta}(p)(\mathcal{U}),F)$, contradicting our assumption.
By Fact 2.5, the binding group $\mathrm{Aut}_{L}(s|_{L}/\mathcal{C})$ is a definable subgroup of $\mathrm{Aut}_{F}(s/\mathcal{C})$. Moreover by Lemma 2.8 the binding group $\mathrm{Aut}_{F}(s/\mathcal{C},\log_{\delta}(p))$ is a normal $F$-definable subgroup of $\mathrm{Aut}_{F}(s/\mathcal{C})$, and, by a proof similar to that of Fact 2.5, is also seen to be a definable subgroup of $\mathrm{Aut}_{L}(s|_{L}/\mathcal{C})$. Since $\mathrm{Aut}_{F}(s/\mathcal{C},\log_{\delta}(p))$ is infinite, as $G_{m}(\mathcal{C})$ is connected and strongly minimal, it must be equal to $\mathrm{Aut}_{L}(s|_{L}/\mathcal{C})$. In particular $\mathrm{Aut}_{L}(s|_{L}/\mathcal{C})$ is normal in $\mathrm{Aut}_{F}(s/\mathcal{C})$.
Note that $\operatorname{RM}(s)=\operatorname{RM}(s|_{L})=\operatorname{RM}(p_{\log_{\delta}(a)})=1$, and as $F$ is algebraically closed, the type $s$ is strongly minimal. Moreover, as $s|_{L}$ is weakly $\mathcal{C}$-orthogonal, forking calculus along with an automorphism argument yields that $s$ is weakly $\mathcal{C}$-orthogonal, and in particular $\mathrm{Aut}_{F}(s/\mathcal{C})$ acts transitively and faithfully on $s$. As it has a normal definable subgroup isomorphic to $G_{m}(\mathcal{C})$, Fact 4.2 implies that $\mathrm{Aut}_{F}(s/\mathcal{C})$ is definably isomorphic to $G_{m}(\mathcal{C})$.